Find an equation in point-slope form of the line having the specified slope and containing the point indicated.
m = -8; (6, 9)

Answer by stanbon(75887)

Find an equation in point-slope form of the line having the specified slope and containing the point indicated.
m = -8; (6, 9)
--------------------------
y = -8x + b and y=9 when x = 6
9 = -8*6 + b
9 = -48 + b
b = 57
---------------
y = -8x + 57
---------
Cheers,
Stan H.